OpenStudy (madgirlwithabluebox):

Use the unit circle to find the inverse function value in degrees. will medal

OpenStudy (madgirlwithabluebox):

\[\cos -1(\frac{ \sqrt{2} }{ 2 })\]

OpenStudy (madgirlwithabluebox):

@peterg2001 thats the wrong one

OpenStudy (peterg2001):

sorry found on google

OpenStudy (madgirlwithabluebox):

its fine thanks anyway :D